Find an SVD of the indicated matrix. (Enter each matrix in the form [[row 1], [row 2], ...), where each row is a comma-separaPPlease be as clear as possible. Thanks
Answer #1

Find Singular Value Decomposition (SVD) of a Matrix ...
[ 5

12 ]

Solution:

A=
5
12


A⋅A′

A⋅A′=
25 60
60 144



Find Eigen vector for A⋅A′

|A⋅A′-λI|=0

(25-λ) 60
60 (144-λ)
= 0



∴(25-λ)×(144-λ)-60×60=0

∴(3600-169λ+λ2)-3600=0

∴(λ2-169λ)=0

∴λ(λ-169)=0

∴λ=0or(λ-169)=0

∴ The eigenvalues of the matrix A⋅A′ are given by λ=0,169,

1. Eigenvectors for λ=169

v1=
0.4167
1


2. Eigenvectors for λ=0

v2=
-2.4
1



For Eigenvector-1 (0.4167,1), Length L = √0.41672+12=1.0833

So, normalizing gives u1=(0.41671.0833,11.0833)=(0.3846,0.9231)

For Eigenvector-2 (-2.4,1), Length L = √(-2.4)2+12=2.6

So, normalizing gives u2=(-2.42.6,12.6)=(-0.9231,0.3846)

A′⋅A

A′⋅A=
[ 169 ]



Find Eigen vector for A′⋅A

|A′⋅A-λI|=0

(169-λ)
= 0



∴(169-λ)=0

∴(-λ+169)=0

∴-(λ-169)=0

∴(λ-169)=0

∴ The eigenvalues of the matrix A′⋅A are given by λ=169,

1. Eigenvectors for λ=169

v1=
[ 1 ]



For Eigenvector-1 (1), Length L = √12=1

So, normalizing gives v1=(11)=(1)

Solution

∴Σ=
√169
0
=
13
0


∴U= [u1,u2] =
0.3846 -0.9231
0.9231 0.3846



V is found using formula vi=1σiAT⋅ui

∴V=
[ 1 ]

----------------------------------------===============================--------------------------------------------------------

Solution:

A=
[ 8 15 ]


A⋅A′

A⋅A′=
[ 289 ]



Find Eigen vector for A⋅A′

|A⋅A′-λI|=0

(289-λ)
= 0



∴(289-λ)=0

∴(-λ+289)=0

∴-(λ-289)=0

∴(λ-289)=0

∴ The eigenvalues of the matrix A⋅A′ are given by λ=289,

1. Eigenvectors for λ=289

v1=
[ 1 ]



For Eigenvector-1 (1), Length L = √12=1

So, normalizing gives u1=(11)=(1)

A′⋅A

A′⋅A=
64 120
120 225



Find Eigen vector for A′⋅A

|A′⋅A-λI|=0

(64-λ) 120
120 (225-λ)
= 0



∴(64-λ)×(225-λ)-120×120=0

∴(14400-289λ+λ2)-14400=0

∴(λ2-289λ)=0

∴λ(λ-289)=0

∴λ=0or(λ-289)=0

∴ The eigenvalues of the matrix A′⋅A are given by λ=0,289,

1. Eigenvectors for λ=289

v1=
0.5333
1


2. Eigenvectors for λ=0

v2=
-1.875
1



For Eigenvector-1 (0.5333,1), Length L = √0.53332+12=1.1333

So, normalizing gives v1=(0.53331.1333,11.1333)=(0.4706,0.8824)

For Eigenvector-2 (-1.875,1), Length L = √(-1.875)2+12=2.125

So, normalizing gives v2=(-1.8752.125,12.125)=(-0.8824,0.4706)

Solution

∴Σ=
[ √289 0 ]
=
[ 17 0 ]


∴U= [u1] =
[ 1 ]



V is found using formula vi=1σiAT⋅ui

∴V=
0.4706 -0.8824
0.8824 0.4706

-----------------------------------------------==========================------------------------------------------------------------

Find Singular Value Decomposition (SVD) of a Matrix ...
[0 1

1 0

2- 2]

Solution:

A=
0 1
1 0
2 -2


A⋅A′

A⋅A′=
1 0 -2
0 1 2
-2 2 8



Find Eigen vector for A⋅A′

|A⋅A′-λI|=0

(1-λ) 0 -2
0 (1-λ) 2
-2 2 (8-λ)
= 0



∴(1-λ)((1-λ)×(8-λ)-2×2)-0(0×(8-λ)-2×(-2))+(-2)(0×2-(1-λ)×(-2))=0

∴(1-λ)((8-9λ+λ2)-4)-0(0-(-4))-2(0-(-2+2λ))=0

∴(1-λ)(4-9λ+λ2)-0(4)-2(2-2λ)=0

∴(4-13λ+10λ2-λ3)-0-(4-4λ)=0

∴(-λ3+10λ2-9λ)=0

∴-λ(λ-1)(λ-9)=0

∴λ=0or(λ-1)=0or(λ-9)=0

∴ The eigenvalues of the matrix A⋅A′ are given by λ=0,1,9,

1. Eigenvectors for λ=9

v1=
-0.25
0.25
1


2. Eigenvectors for λ=1

v2=
1
1
0


3. Eigenvectors for λ=0

v3=
2
-2
1



For Eigenvector-1 (-0.25,0.25,1), Length L = √(-0.25)2+0.252+12=1.0607

So, normalizing gives u1=(-0.251.0607,0.251.0607,11.0607)=(-0.2357,0.2357,0.9428)

For Eigenvector-2 (1,1,0), Length L = √12+12+02=1.4142

So, normalizing gives u2=(11.4142,11.4142,01.4142)=(0.7071,0.7071,0)

For Eigenvector-3 (2,-2,1), Length L = √22+(-2)2+12=3

So, normalizing gives u3=(23,-23,13)=(0.6667,-0.6667,0.3333)

A′⋅A

A′⋅A=
5 -4
-4 5



Find Eigen vector for A′⋅A

|A′⋅A-λI|=0

(5-λ) -4
-4 (5-λ)
= 0



∴(5-λ)×(5-λ)-(-4)×(-4)=0

∴(25-10λ+λ2)-16=0

∴(λ2-10λ+9)=0

∴(λ-1)(λ-9)=0

∴(λ-1)=0or(λ-9)=0

∴ The eigenvalues of the matrix A′⋅A are given by λ=1,9,

1. Eigenvectors for λ=9

v1=
-1
1


2. Eigenvectors for λ=1

v2=
1
1



For Eigenvector-1 (-1,1), Length L = √(-1)2+12=1.4142

So, normalizing gives v1=(-11.4142,11.4142)=(-0.7071,0.7071)

For Eigenvector-2 (1,1), Length L = √12+12=1.4142

So, normalizing gives v2=(11.4142,11.4142)=(0.7071,0.7071)

Solution

∴Σ=
√9 0
0 √1
0 0
=
3 0
0 1
0 0


∴U= [u1,u2,u3] =
-0.2357 0.7071 0.6667
0.2357 0.7071 -0.6667
0.9428 0 0.3333



V is found using formula vi=1σiAT⋅ui

∴V=
0.7071 0.7071
-0.7071 0.7071